The mass of a certain planet P is one-hundredth that of the earth while the radius is one quarter of the earth's radius. If the acceleration due to gravity on the earth is 10m/s2,what is the value of P?
1
Expert's answer
2019-09-30T10:05:59-0400

The acceleration due to gravity


"g=G\frac{M}{R^2}"

So

"\frac{g_P}{g_E}=\frac{M_P}{M_E}\times \frac{R_E^2}{R_P^2}"

"\frac{g_P}{10}=\frac{1}{100}\times 16"

"g_P=1.6\:\rm m/s^2"